OpenStudy (anonymous):

Medal Solve by determinants. x + y + z = 21 x=y +z = 3 x +y - z = 1

jimthompson5910 (jim_thompson5910):

is there a typo in the second equation `x=y +z = 3` ??

jimthompson5910 (jim_thompson5910):

there are 2 equal signs there

OpenStudy (anonymous):

yes sorry its x - y + z = 3

jimthompson5910 (jim_thompson5910):

ok thanks

jimthompson5910 (jim_thompson5910):

were you able to set up the matrix?

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

what matrix did you get

OpenStudy (anonymous):

1 1 1 1 -1 1 1 1 -1

OpenStudy (anonymous):

m i right?

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

that's your coefficient matrix

jimthompson5910 (jim_thompson5910):

do you know how to take the determinant of that matrix?

OpenStudy (anonymous):

no

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

focus on the first two columns